OpenStudy (anonymous):

simplify the radical below: √18x^7

OpenStudy (cwrw238):

√18 = √9 * √2 = 3√2

OpenStudy (anonymous):

cou;d u help me on 1 more problem?

OpenStudy (cwrw238):

ii am assuming the square root refers to the 18 only if your answer is 3√2x^7

OpenStudy (cwrw238):

ok

OpenStudy (anonymous):

oh i think there is more to do

OpenStudy (anonymous):

\[\sqrt{x^7}=\sqrt{x^6\times x}=x^3\sqrt{x}\]

OpenStudy (anonymous):

yeah